OpenStudy (anonymous):

which of the following pairs of lines are neither parallel nor perpendicular? a. x-2y=7 and 3x-6y=13 b. x-2y=7 and 2x+y =9 c. x-2y =7 and 2x-4y =5 d.x-2y=7 and x+2y =7

OpenStudy (whpalmer4):

If the lines are parallel, their slopes will be identical. If the lines are perpendicular, the product of their slopes will be -1. Rearrange each equation into slope-intercept form: \(y=mx+b\) where \(m\) is the slope. You can compare the slopes after doing that.

OpenStudy (anonymous):

@whpalmer4 is the answer is b?

OpenStudy (whpalmer4):

What leads you to that conclusion?

OpenStudy (whpalmer4):

What are the slopes you got for the lines x-2y=7 and 2x+y=9?